# Instituto de Física Teórica UAM-CSIC, Universidad Autónoma de Madrid, C/ Nicolás Cabrera 13-15, 28049 Madrid, Spain Departamento de Física Teórica, Universidad Autónoma de Madrid, M-15, 28049 Madrid, Spain Simulating Electromagnetic Cascades with Lorentz Invariance Violation

### v1

@@ 1 Introduction
QG theories attempt to describe the four fundamental forces within a single theoretical framework [4]. It generally strives to bridge the gap between the quantum field theory approach of the SM and the differential geometry framework employed in general relativity (GR).
@@
There have been several works positing that QG lead to a phenomenon known as Lorentz invariance violation (LIV) [10, 11, 12, 13, 14]. Lorentz symmetry can be broken if at least one of the pillars upon which it rests is broken, namely the principle of relativity, which ensures the equivalence of the laws of physics in all inertial frames, or the constancy of the speed of light. The immediate phenomenological consequence of LIV is, thus, a modification of the energy-momentum dispersion relation and spatio-temporal changes. Note that some QG frameworks also feature modified dispersion relations whilst merely deforming Lorentz symmetry. This, for example, is the case for deformed (or doubly) special relativity (DSR) [15, 16].
@@
In this work we focus on the expected signatures of LIV on gamma rays from cosmological sources, considering their intergalactic propagation [27]. During their journey from an astrophysical object to Earth, sufficiently energetic photons from an astrophysical object like a blazar or a gamma-ray burst interact with radiation pervading the universe, such as the cosmic microwave background (CMB) and the extragalactic background light (EBL), producing an electron-positron pair. These pairs, in turn, interact with background photons via inverse Compton scattering, transferring a substantial fraction of their energy to the photons resulting from it. These photons, depending on their energy, can trigger these interactions yet again, until the high-energy photons no longer have sufficient energy to start the reaction. This energy threshold is modified if LIV is acting [28, 22], changing the interaction length for pair production and igniting or quenching the cascade.
@@ 2 Particle Propagation with Modified Dispersion Relations
The consequences of these modifications are two-fold [22]. Firstly, the threshold of a given reaction is altered, leading to a modification in the associated propagation length, as described in section 2.1. Secondly, the presence of LIV enables new reactions that would otherwise not be possible within the confines of Lorentz invariance. Examples of such reactions include the spontaneous decay of photons into pairs or photons, as well as the occurrence of the vacuum Cherenkov effect for electrons, both described in section 2.2.
@@ 2.2 New Electromagnetic Processes
Within our framework, we make the assumption that if the parent’s energy exceeds the threshold for vacuum Cherekov (VC), it will emit one or more photons until its energy falls below this threshold. To streamline our approach, we adopt the same methodology as we did for photons: if the electron or positron’s energy is above the threshold, we reduce its energy to match the threshold value. Conversely, if the energy falls below the threshold, we do not consider the VC effect. This evaluation is conducted at each propagation step. The specific threshold value we utilize is derived from the findings presented in [22], which we summarize in the following section.
@@
The second “new” reaction, photon decay, can be described by
This process involves the spontaneous transformation of a high-energy photon into an electron-positron pair without interaction with a background photon. Similarly to the VC mechanism, we assume that when the energy of the photon is above a certain threshold for photon decay, it will continually decay into electron-positron pairs until its energy drops below this threshold. Consequently, we employ a simplified “binary” approach: if the photon’s energy is above the threshold, it is reduced to the threshold value, while if the energy is below the threshold, photon decay does not occur. We compute the specific threshold value based on the results presented in [22], which we summarize in the next paragraphs.
We can now determine the minimum threshold energy that a particle has to have in order for VC to take place. This is shown in figure 3.
@@
The threshold energy for photon decay (PD) is shown in figure 4.
@@ 3 Simulations of Electromagnetic Cascades with Lorentz Invariance Violation
To simulate the propagation of gamma-ray–induced electromagnetic cascades, we extend the CRPropa code [31, 32], tapping into the already existing structure of this framework to calculate the relevant kinematics in the presence of LIV [33]. We also implemented new modules for treating vacuum Cherekov and photon decay.
In the following sections we discuss the phenomenology of LIV using simulations obtained with these tools. We select a few scenarios to motivate the discussion, but it is not our goal to compare simulations with observations at this stage.
We start off by considering the case of LIV affecting the propagation of photons (section 3.1). Then, in section 3.2, we discuss the case of electrons. In section 3.3 we provide comprehensive simulations comprising all effects considered, namely pair production (PP), inverse Compton scattering (ICS), PD, and VC. In section 3.4 we examine the effect of intrinsic source properties on gamma-ray fluxes within the context of LIV.
@@ 3.1 Simulations: Photon Propagation
Due to absorption by background cosmological radiation fields, namely the CMB and the EBL, a strong suppression of the flux measured at Earth is expected, owing to pair-production interactions. In addition, LIV generally allows for photon decay as well. These two effects, combined, affect the spectrum. This is shown in figure 5.
@@ 3.2 Simulations: Electron Propagation
We proceed to examine the impact of vacuum Cherekov on the propagation dynamics. To achieve this, we investigate a scenario similar to that for photons, but this time considering ICS instead of PP, and VC in place of PD. The results are presented in figure 6.
@@ 3.3 Simulations: Full LIV Phenomenology
Having isolated each individual effect in the simulations, we now proceed to study a scenario with all relevant phenomenology. We compare the special-relativistic case with the ones including LIV. The results are show in figure 7.
The superluminal (purple shades) and subluminal (orange shades) cases of Fig. 7 exhibit very different phenomenological signatures. The former, in particular, can be understood as an increased transparency of the universe, thus explaining why a signal would be seen at energies above the expected suppression due to the EBL absorption of photons.
@@ 3.4 Simulations: Effects of the Intrinsic Properties of the Source
where denotes the spectral index. We analyse a couple of cases for the same two sources used in section 3.1. The results are shown in figures 8 and 9.
@@ 5 Conclusions and Outlook
In the future, we envision expanding our analysis in several directions. One avenue of exploration involves refining our treatments of photon decay and vacuum Cherekov, which currently are done only for LIV of order and . We also intend to generalise the LIV formalism to allow superposition of order (multiple values of acting together). Finally, we intend to include effects of magnetic fields to understand how they can affect the charged leptonic component of electromagnetic cascades.

### v2

# 3Instituto de Física Teórica UAM-CSIC, Universidad Autónoma de Madrid, C/ Nicolás Cabrera 13-15, 28049 Madrid, Spain 4Departamento de Física Teórica, Universidad Autónoma de Madrid, M-15, 28049 Madrid, Spain 5Sorbonne Université, CNRS, UMR 7095, Institut d’Astrophysique de Paris, 98 bis bd Arago, 75014 Paris, France

@@ 1 Introduction
QG theories attempt to describe the four fundamental forces within a single theoretical framework [4]. It generally strives to bridge the gap between the quantum field theory approach of the SM and the differential geometry framework employed in general relativity (GR). This task presents a significant challenge due to the classical nature of GR, which inherently complicates its combination with any form of quantization.
@@
There have been several works positing that QG lead to a phenomenon known as Lorentz invariance violation [10, 11, 12, 13, 14]. Lorentz symmetry can be broken if at least one of the pillars upon which it rests is broken, namely the principle of relativity, which ensures the equivalence of the laws of physics in all inertial frames, or the constancy of the speed of light. The immediate phenomenological consequence of LIV is, thus, a modification of the energy-momentum dispersion relation and spatio-temporal changes. Note that some QG frameworks also feature modified dispersion relations whilst merely deforming Lorentz symmetry. This, for example, is the case for deformed (or doubly) special relativity (DSR) [15, 16].
@@
In this work we focus on the expected signatures of LIV on gamma rays from cosmological sources, considering their intergalactic propagation [27]. During their journey from an astrophysical object to Earth, sufficiently energetic photons from an astrophysical object like a blazar or a gamma-ray burst interact with radiation pervading the universe, such as the cosmic microwave background (CMB) and the extragalactic background light (EBL), producing an electron-positron pair. These pairs, in turn, interact with background photons via inverse Compton scattering, transferring a substantial fraction of their energy to the photons resulting from it. These photons, depending on their energy, can trigger these interactions yet again, until the high-energy photons no longer have sufficient energy to start the reaction. Therefore, a single primary photon of sufficiently high energy can result in copious amounts of secondaries, tertiaries, and higher-order gamma rays. This energy threshold is modified if LIV is acting [28, 22], changing the interaction length for pair production and igniting or quenching the cascade.
@@ 2 Particle Propagation with Modified Dispersion Relations
The consequences of these modifications are two-fold [22]. Firstly, the threshold of a given reaction is altered, leading to a modification in the associated propagation length, as described in Sec. 2.1. Secondly, the presence of LIV enables new reactions that would otherwise not be possible within the confines of Lorentz invariance. Examples of such reactions include the spontaneous decay of photons into pairs or photons, as well as the occurrence of the vacuum Cherenkov effect for electrons, both described in Sec. 2.2.
@@ 2.2 New Electromagnetic Processes
Within our framework, we make the assumption that if the parent’s energy exceeds the threshold for vacuum Cherekov (VC), it will emit one or more photons until its energy falls below this threshold. To streamline our approach, we adopt a “binary” methodology: if the electron or positron’s energy is above the threshold, we reduce its energy to match the threshold value. Conversely, if the energy falls below the threshold, we do not consider the VC effect. This evaluation is conducted at each propagation step. The specific threshold value we utilize is derived from the findings presented in [22], which we summarize in the following section.
@@
The second “new” reaction, photon decay (PD), can be described by
This process involves the spontaneous transformation of a high-energy photon into an electron-positron pair without interaction with a background photon. Similarly to the VC mechanism, we assume that when the energy of the photon is above a certain threshold for photon decay, it will continually decay into electron-positron pairs until its energy drops below this threshold. Consequently, we again employ a simplified “binary” approach: if the photon’s energy is above the threshold, it is reduced to the threshold value, while if the energy is below the threshold, photon decay does not occur. This effectively corresponds to a hard cut-off in the photon spectrum, which is a standard approach in this context in the literature [31, 32, 33]. We compute the specific threshold value based on the results presented in [22], which we summarize in the next paragraphs.
We can now determine the minimum threshold energy that a particle has to have in order for VC to take place. This is shown in Fig. 3.
It follows from Fig. 3 that, for the superluminal case (left panel), the value of completely determines the energy threshold of VC regardless of . For the subluminal case (right panel) the situation is reversed, with exerting little influence on the energy thresholds, which decrease proportionally to . Note that the subluminal case does not allow for the region of the parameter space left blank, according to Eq. (14).
@@
The threshold energy for PD is shown in Fig. 4.
Fig. 4 indicates that the energy threshold for photon decay in the superluminal case (left panel) decreases as and jointly increase. The subluminal case (right panel) has a forbidden region, according to Eq. (17), wherein photon decay does not take place. Interestingly, within the allowed region of the parameter space, the energy threshold decreases for large values of , as expected, but it decreases for lower .
@@ 3 Simulations of Electromagnetic Cascades with Lorentz Invariance Violation
To simulate the propagation of gamma-ray–induced electromagnetic cascades, we extend the CRPropa code [34, 35], tapping into the already existing structure of this framework to calculate the relevant kinematics in the presence of LIV [36].
CRPropa111https://crpropa.desy.de/ is an advanced framework for simulating the propagation of high and ultra-high-energy messengers. It implements the main interactions relevant for the propagation of each type of messenger, using a Monte Carlo approach. For this work, we modified the existing modules that handle electromagnetic interactions [37], which compute the kinematics of pair production and inverse Compton scattering. These modules make use of the routines provided in CRPropa’s auxiliary repository 222https://github.com/CRPropa/CRPropa3-data. We also implemented new modules for treating vacuum Cherekov and photon decay, taking advantage of the code’s modular structure.
In the following sections we discuss the phenomenology of LIV using simulations obtained with these tools. We select a few scenarios to motivate the discussion, but it is not our goal to compare simulations with observations at this stage. Note that we restrict our analysis to , considering the substantial computational costs of performing Monte Carlo simulations at higher energies. For this reason, we only present results for the case, since for there is not significant different between the Lorentz invariant and the LIV scenarios in the energy range to which we are limited.
We start off by considering the case of LIV affecting the propagation of photons (Sec. 3.1). Then, in Sec. 3.2, we discuss the case of electrons. In Sec. 3.3 we provide comprehensive simulations comprising all effects considered, namely pair production (PP), inverse Compton scattering (ICS), PD, and VC. In Sec. 3.4 we examine the effect of intrinsic source properties on gamma-ray fluxes within the context of LIV.
@@ 3.1 Simulations: Photon Propagation
Due to absorption by background cosmological radiation fields, namely the CMB and the EBL, a strong suppression of the flux measured at Earth is expected due to pair-production interactions, at some energy range, depending on the source distance, according to Fig. 1. For instance, any source more distant than cannot contribute to the gamma-ray flux measured at Earth around . In addition, LIV generally allows for photon decay as well. These two effects, combined, affect the spectrum. This is shown in Fig. 5.
@@ 3.2 Simulations: Electron Propagation
We proceed to examine the impact of vacuum Cherekov on the propagation dynamics. To achieve this, we investigate a scenario similar to that for photons, but this time considering ICS instead of PP, and VC in place of PD. The results are presented in Fig. 6.
@@ 3.3 Simulations: Full LIV Phenomenology
Having isolated each individual effect in the simulations, we now proceed to study a scenario with all relevant phenomenology. We compare the special-relativistic case with the ones including LIV. The results are show in Fig. 7.
The superluminal (purple shades) and subluminal (orange shades) cases of Fig. 7 exhibit very different phenomenological signatures. The latter, in particular, is a clear manifestation of an increased transparency of the universe, thus explaining why a signal is seen at energies above the expected suppression due to the EBL absorption of photons.
@@ 3.4 Simulations: Effects of the Intrinsic Properties of the Source
where denotes the spectral index. We analyse a couple of cases for the same two sources used in Sec. 3.1. The results are shown in figures 8 and 9.
@@ 4 Discussion
It should be noted here that the aforementioned “binary” treatment of these new reactions (in this paper specifically photon decay and the vacuum Cherenkov effect) is a very simplified approach. Using a more sophisticated treatment, taking into account the full spectrum of the corresponding secondaries, this might change the overall observed spectrum. However, we chose the simplification, on the one hand, due to the fact that a full treatment involves non-trivial quantum-electrodynamical calculations which are beyond the scope of the phenomenological approach of the present paper. On the other hand, VC and PD usually occur on length scales much shorter than ICS and PP, such that the particles undergoing the latter two reactions are already the result of the two former reactions, and therefore the effects may, at least to some degree as a compromise, be treated separately.
@@
Other factors would affect the results presented. For instance, the EBL model determines the interaction rates of pair production (see Fig. 1). The effect of the EBL in the presence of LIV would not qualitatevely differ from the Lorentz-invariant case.
IGMF could affect the charged component of the cascade, potentially altering the energy spectra expected at Earth [27]. In fact, this could greatly impact gamma-ray–based estimates of intergalactic magnetic fields (e.g., [46, 47, 48, 49, 50]). Similarly, the interplay between the geometry of emission and IGMFs could change spectral and angular signatures of gamma rays from misaligned jetted sources [51].
We have performed this investigation only for electromagnetic interactions, but analogous signatures are expected for other types of travelling messengers, like UHECRs. In particular, the GZK effect due to the interaction of energetic protons with the CMB, if it occurs, could be substantially altered in the presence of LIV, as studied in [52, 53, 54, 55].
@@ 5 Conclusions and Outlook
In the future, we envision expanding our analysis in several directions. One avenue of exploration involves refining our treatments of photon decay and vacuum Cherekov, which currently are done only for LIV of order and , in particular by using a complete quantumelectrodynamic treatment, such as, for example, the one presented in [56] for . We also intend to generalise the LIV formalism to allow superposition of order (multiple values of acting together). Finally, we intend to include effects of magnetic fields to understand how they can affect the charged leptonic component of electromagnetic cascades.
@@
Data Availability Statement
The code used for the simulations in this manuscript is being prepared to be made public. In the meantime, it can be made available upon reasonable request. The data that support the findings of this study are available upon reasonable request from the authors.